OpenStudy (anonymous):

Identify a reflection about the x-axis or a reflection about the y-axis of ƒ(x) = x2 + 4x3 by observing the equation of the function g(x) = x2 – 4x3

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=x^2+4x^3 }\) \(\large\color{blue}{ g(x)=x^2-4x^3 }\) sure this is the information? maybe one sign is missing?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

no thats all

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=x^2+4x^3 \\ g(x) = -x^2 -4x^3 }\) this is when the g(x) is a reflection of f(x) across the x-axis.

OpenStudy (solomonzelman):

Not when \(\large\color{blue}{ g(x) = x^2 -4x^3 }\)